OpenStudy (anonymous):

find all exact solutions in [0,2pi) of the trigonometric equation 1-2cos^2(x/2) = sin^2(x)

OpenStudy (loser66):

your idea?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

\[1-2\cos ^{2}(\frac{ x }{ 2 }) = \sin ^{2}(x) \in [0,2\pi)\]

OpenStudy (loser66):

I think you have some strategy to approach the problem, right? give me your idea. I have to know where you stuck

OpenStudy (anonymous):

To be honest, I'm completely lost. Have no clue how to start

OpenStudy (loser66):

how about double angle rule? nothing?

OpenStudy (anonymous):

double angle rule = a little

OpenStudy (anonymous):

\[\sin 2u = 2 \sin u \cos u\]

OpenStudy (loser66):

ok, \[cos \dfrac{x}{2}= \pm\sqrt{\dfrac{1+cos x}{2}}\] so that \( cos^2(x/2) =\dfrac{1+cos x}{2}\) --> \(2cos^2(x/2)= 1+cosx\)

OpenStudy (anonymous):